# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### 20 Jan 2024 v1

@@ Appendix C TAPIS performance in GraSP
Figures 26, 27, and 28 depict qualitative segmentation results from TAPIS on the GraSP dataset. As mentioned previously, TAPIS can perform accurate segmentation of complex images. In addition, Figure 26 shows that TAPIS can segment instances of varying sizes, including bigger instrument arms and small tips. We show examples of inaccurate predictions in Figure 28. We conduct an error mode analysis and observe cases where two instrument instances are predicted as one, given a high degree of overlap between them (see first row, upper left corner). Second, we observe cases where an instrument is classified as another, given a problematic view of the instrument. For example, the fifth row confuses Bipolar Forceps with Prograsp Forceps. These instruments are similar and difficult to distinguish from a profile view. Third, we observe cases where it wrongly identifies other objects as instruments. For instance, the third row identifies an instrument for recollecting waste as a Laparoscopic Instrument.
@@
Phase and step recognition
On the one hand, Figures 29 and 30 present qualitative outcomes of TAPIS in the phase recognition task. Notably, our model encounters challenges in recognizing phases with shorter durations. In instances where recognition occurs, the model tends to identify longer time segments than their actual duration. Conversely, phases with longer durations exhibit more accurate recognition by the model.
On the other hand, Figures 31 and 32 illustrate results in the steps recognition task. Notably, the model demonstrates noise in recognizing steps with longer durations while performing well in predicting short steps. Additionally, timestamps may experience shifts, indicating challenges in defining transitions between steps.
Action detection

### 26 Jan 2024 v2

@@ Appendix D TAPIS performance in GraSP
Figures 15, 16, and 17 depict qualitative segmentation results from TAPIS on the GraSP dataset. As mentioned previously, TAPIS can perform accurate segmentation of complex images. In addition, Figure 15 shows that TAPIS can segment instances of varying sizes, including bigger instrument arms and small tips. We show examples of inaccurate predictions in Figure 17. We conduct an error mode analysis and observe cases where two instrument instances are predicted as one, given a high degree of overlap between them (see first row, upper left corner). Second, we observe cases where an instrument is classified as another, given a problematic view of the instrument. For example, the fifth row confuses Bipolar Forceps with Prograsp Forceps. These instruments are similar and difficult to distinguish from a profile view. Third, we observe cases where it wrongly identifies other objects as instruments. For instance, the third row identifies an instrument for recollecting waste as a Laparoscopic Instrument. Additionally, Tables 7 and 8 show the per-class results on instrument segmentation explained in the Instrument Segmentation Experiments Section.
Phase and step recognition
On the one hand, Figures 18 and 19 present qualitative outcomes of TAPIS in the phase recognition task. Notably, our model encounters challenges in recognizing phases with shorter durations. In instances where recognition occurs, the model tends to identify longer time segments than their actual duration. Conversely, phases with longer durations exhibit more accurate recognition by the model.
On the other hand, Figures 20 and 21 illustrate results in the steps recognition task. Notably, the model demonstrates noise in recognizing steps with longer durations while performing well in predicting short steps. Additionally, timestamps may experience shifts, indicating challenges in defining transitions between steps. Tables 9 and 10 present the per-class quantitative results on phases and steps recognition as explained in the Phases and Steps Experiments Section.
Action detection